XYZ Corporation has the following financial information for the previous year: 


Sales: $8M, Profit Margin (PM)-8%, Current Asset (CA) = $2M, Fixed Asset (FA) = $6M, NWC-SIM, LTD = $3M 


Compute the ROE using the DuPont Analysis.
Answer #1

ROE = Net Profit Margin x Asset Turnover x Financial Leverage = Net Income / Sales x Sales / Assets x Assets / Equity = Net Income / Equity

Here, NPM = 8%

Assets = CA + FA = $8M

Asset Turnover = Sales / Asset = 8 / 8 = 1.0

Total Liabilities = CL + LTD = (CA - NWC) + LTD = 1 + 3 = 4

=> Equity = Assets - Liabilities = 8 - 4 = $4M

Financial Leverage = Asset / Equity = 8 / 4 = 2.0

=> ROE = 8% x 1.0 x 2.0 = 16.0%


answered by: gavin